1)   An instruction is stored at location 300 with its address field at location 301.

          The address field has the value 400. A processor register R 1 contains the

           number 200. Evaluate the effective address if the addressing mode of the

            instruction is (a) direct; (b) immediate; (c) relative; (d) register indirect;

                   (e) index with R1 as the index register.  


1
Expert's answer
2021-11-19T01:54:36-0500

The description given in the assignment can be represented like this:

Location _ Contents

300 _ opcode ;the instruction operation code

301 _ 400 ;address field of the above instruction


(a) direct addressing

Direct addressing means that the address field contains the address of memory location the instruction is supposed to work with (where an operand "resides").

Effective address would therefore be 400.


(b) immediate addressing

Immediate addressing means that the address field contains the operand itself.

Effective address would therefore be 301.


(c) relative addressing

Relative addressing means that the address field contains offset to be added to the program counter to address a memory location of the operand.

Effective address would therefore be 301 + 400 = 701.


(d) register indirect addressing

Register indirect addressing means that the address of an operand is in the register. The address field in this case contains just another operand.

Effective address would therefore be in R1 = 200.


(e) indexed addressing with R1 as index register

There are several possible indexed addressing modes but in this case (tehre is an address field) it is co called "indexed absolute" addressing.

In indexed absolute addressing the effective address is calculated by taking the contnts of the address field and adding the contents of the index register.

Effective address would therefore be 400 + R1 = 400 + 200 = 600